# Primal-Dual iLQR for GPU-Accelerated Learning and Control in Legged Robots

### v1

@@ II-D Parallel Associative Scan
In this section, we will briefly recap how parallel associative scans work and how they can be used to solve the primal problem as shown for the first time by [18]. Given a sequence of elements $a_{1},\dots,a_{N}$ and an associative operator $\otimes$ defined on them, a parallel associative scan can be used to evaluate the values $s_{1},\dots,s_{N}$ as:
$s_{1}$ $=a_{1},$
$s_{2}$ $=a_{1}\otimes a_{2},$
$\vdots$
$s_{N}$ $=a_{1}\otimes a_{2}\otimes\cdots\otimes a_{N}.$
in $\mathcal{O}(\log{T})$ as shown by [33]. Thanks to the associative property of the $\otimes$ operator we can divide the computation into independent subproblems that are solvable in parallel. To apply such an algorithm to the primal problem solution shown in Sec. II-C we first have to define the Conditional Value Function $V_{i\rightarrow{}j}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ as the cost related to the optimal trajectory that goes from the state $\boldsymbol{x}_{i}$ at node $i$ in the horizon to state $\boldsymbol{x}_{j}$ at node $j$. For the problem in (II-B) we can write the Conditional Value Function as the $\min\limits_{u_{i:j}}\sum_{i}^{j}{\boldsymbol{l}_{i}(\boldsymbol{x}_{i}\boldsymbol{u}_{i})}$subject to the dynamics in (1b). We can define the Lagrangian multiplier $\boldsymbol{\eta}$ such that:
@@
where the tilde $\tilde{\cdot}$ is used to distinguish the values related to the conditional value function from the one in (4), while the subscripts are used to specify the initial and final state considered. Särkkä et al. [18] proved that given a third state $\boldsymbol{x}_{k}$ and given that $V_{i\rightarrow{k}}$ and $V_{k\rightarrow{j}}$ are in the same form of (7), $V_{i\rightarrow{j}}$ will still be in the form of (7) and the Lagrangian is characterized by such combination rule:
@@
We now use the conditional value function $V_{i\rightarrow{i+1}}$ defined between two subsequent state $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{i+1}$ to initialize the values $a_{i}$ of a reverse associative scan. $V_{i\rightarrow{i+1}}$ for $i=0,\ldots,N-1$ is characterized by:
$\boldsymbol{\tilde{P}}_{i,i+1}$ $=\boldsymbol{Q_{i}}-\boldsymbol{S_{i}}\boldsymbol{R}_{i}^{-1}\boldsymbol{S}_{i}^{\top},$ (9)
$\boldsymbol{\tilde{p}}_{i,i+1}$ $=\boldsymbol{q_{i}}-\boldsymbol{S_{i}}\boldsymbol{R}_{i}^{-1}r_{i},$
$\boldsymbol{\tilde{A}}_{i,i+1}$ $=\boldsymbol{A_{i}}-\boldsymbol{B_{i}}\boldsymbol{R}_{i}^{-1}S_{i}^{\top},$
$\boldsymbol{\tilde{C}}_{i,i+1}$ $=\boldsymbol{B_{i}}\boldsymbol{R_{i}}^{-1}\boldsymbol{B}_{i}^{\top},$
@@
$\boldsymbol{\tilde{P}}_{N,N+1}$ $=\boldsymbol{Q}_{N+1},$ (10)
$\boldsymbol{\tilde{p}}_{N,N+1}$ $=\boldsymbol{q}_{N+1},$
$\boldsymbol{\tilde{A}}_{N,N+1}$ $=0$
$\boldsymbol{\tilde{C}}_{N,N+1}$ $=0$
$\boldsymbol{\tilde{b}}_{N,N+1}$ $=0$
From the solution of the parallel associative scan using the combination rule in (8), we obtain
$V_{i\rightarrow N+1}(\boldsymbol{x}_{i},\boldsymbol{x}_{N+1})=a_{i}\otimes\dots\otimes a_{N+1}.$
Finally, it can be proven that $V_{i\rightarrow{N+1}}(\boldsymbol{x}_{i},\boldsymbol{x}_{N+1})=V(\boldsymbol{x}_{i})$ with $V(\boldsymbol{x}_{i})$ the value function of the problem in (II-B) that is in the form $V(\boldsymbol{x}_{i})=\frac{1}{2}\boldsymbol{x}_{i}^{\top}\boldsymbol{P}_{i}\boldsymbol{x}_{i}-\boldsymbol{p}_{i}^{\top}\boldsymbol{x}_{i}$ and so $\boldsymbol{P}_{i}=\boldsymbol{\tilde{P}}_{i,N+1}$ and $\boldsymbol{\tilde{p}}_{i}=\boldsymbol{p}_{i,N+1}$. Using the parallel associative scan we obtained all $\boldsymbol{P}_{i}$ and $\boldsymbol{p}_{i}$ for $i=0,\,\dots\,,n+1$. We can then use (II-C) to calculate $\boldsymbol{K}$ and $\boldsymbol{k}$ in parallel.
Thanks to the parallel scan implementation on GPU available in JAX, we can get the control policy along the entire horizon in just one pass with an overall computational complexity of $\mathcal{O}(n\log{N}+m)$ as described in [18] in Sec. V-B.
The parallel associative scan can also be used to retrieve the optimal trajectory in $\mathcal{O}(\log{n}\log{N}+\log{m})$. We start plugging into the linearized dynamics the optimal control law we just evaluated, obtaining:
$\boldsymbol{\delta x}_{i+1}=\boldsymbol{\bar{A}}_{i}\boldsymbol{\delta x}_{i}+\boldsymbol{\tilde{b}}_{i}$ (11)
where:
$\boldsymbol{\bar{A}}_{i}=\boldsymbol{A}_{i}+\boldsymbol{B}_{i}\boldsymbol{K}_{i}\quad\text{and}\quad\boldsymbol{\tilde{b}}_{i}=\boldsymbol{B}_{i}\boldsymbol{k}_{i}+\boldsymbol{b}_{i}$
We can define the conditional optimal trajectory as follows:
$\boldsymbol{h}_{i\rightarrow{j}}(\boldsymbol{\delta x}_{i},\boldsymbol{\delta x}_{j})=\boldsymbol{\bar{A}}_{i,j}\boldsymbol{\delta x}_{i}+\boldsymbol{\bar{b}}_{i,j}$ (12)
[18] showed that from the combination of two conditional optimal trajectory $\boldsymbol{h}_{i\rightarrow{k}}$ and $\boldsymbol{h}_{k\rightarrow{j}}$ we obtain $\boldsymbol{h}_{i\rightarrow{j}}$ that can be written in the same form as (12) with:
$\boldsymbol{\bar{A}}_{i,j}=\boldsymbol{\bar{A}}_{i,k}\boldsymbol{\bar{A}}_{k,j}$ (13)
$\boldsymbol{\bar{b}}_{i,j}=\boldsymbol{\bar{A}}_{k,j}\boldsymbol{\bar{b}}_{i,k}+\boldsymbol{\bar{b}}_{k,j}$
@@
$\boldsymbol{\bar{A}}_{i,i+1}=\boldsymbol{\bar{A}}_{i}$ (14)
$\boldsymbol{\bar{b}}_{i,i+1}=\boldsymbol{\bar{b}}_{i}$
@@ II-E Parallel Line Search
$\boldsymbol{x}^{k+1}=\boldsymbol{x}^{k}+\alpha\boldsymbol{\delta x}$ (16)
@@
$\boldsymbol{\lambda}^{k+1}=\boldsymbol{\lambda}^{k}+\alpha\boldsymbol{\delta\lambda}$
The SQP linear rollout allows for the computation to be easily performed in parallel for the horizon. To evaluate $\alpha$, we implemented a filter line search similar to the one implemented in [34]. The filter line search ensures that the step taken in the iteration reduces the constraint violation or the cost. Such a method has already been successfully deployed for legged robot control in [2]. We measure the constraint satisfaction as:
@@
As for [2], when the constraint violation $\theta$ is higher than a threshold, we reject the step if it further increases $\theta$. If the violation is below the threshold and the current step is a descent direction, then we apply an Armijo condition [35]. On the other hand, if the current iterate is not in a descent direction, we require that at least the cost or $\theta$ is decreased to accept the step. To exploit the parallelization on the GPU, we are performing the line search on a fixed set of alpha values at the same time and picking the highest alpha that satisfies the line search conditions.
@@ II-F Practical Implementation
$\boldsymbol{Q}_{N+1}$ $=\nabla_{x}\boldsymbol{\epsilon}_{N+1}^{\top}W_{T}\nabla_{x}\boldsymbol{\epsilon}_{N+1}$ (18)
$\boldsymbol{Q}_{i}$ $=\nabla_{x}\boldsymbol{\epsilon}_{i}^{\top}W_{i}\nabla_{x}\boldsymbol{\epsilon}_{i}$
$\boldsymbol{S}_{i}$ $=\nabla_{x}\boldsymbol{\epsilon}_{i}^{\top}W_{i}\nabla_{u}\boldsymbol{\epsilon}_{i}$
$\boldsymbol{R}_{i}$ $=\nabla_{u}\boldsymbol{\epsilon}_{i}^{\top}W_{i}\nabla_{u}\boldsymbol{\epsilon}_{i}$
@@
$\mathcal{B}(\bar{h},\mu,\delta)\begin{cases}-\mu\ln(\bar{h}),&\text{if }\bar{h}\geq\delta,\\ \frac{\mu}{2}\left(\left(\frac{\bar{h}-2\delta}{\delta}\right)^{2}-1\right)-\mu\ln(\delta),&\text{if }\bar{h}<\delta.\end{cases}$ (19)
where $\bar{h}(\boldsymbol{x},\boldsymbol{u})<0$ is the considered constraint, and $\mu$ and $\delta$ are parameters tuned for each constraint. Following the example of [2], to maintain the convexity of II-B the hessian approximation of the constraints is constructed as
$\boldsymbol{Q}_{N+1}$ $=\nabla_{x}\boldsymbol{\bar{h}}_{N+1}^{\top}\nabla_{\bar{h}\bar{h}}\mathcal{B}_{T}\nabla_{x}\boldsymbol{\bar{h}}_{N+1}$ (20)
$\boldsymbol{Q}_{i}$ $=\nabla_{x}\boldsymbol{\bar{h}}_{i}^{\top}\nabla_{\bar{h}\bar{h}}\mathcal{B}_{i}\nabla_{x}\boldsymbol{\bar{h}}_{i}$
$\boldsymbol{S}_{i}$ $=\nabla_{x}\boldsymbol{\bar{h}}_{i}^{\top}\nabla_{\bar{h}\bar{h}}\mathcal{B}_{i}\nabla_{u}\boldsymbol{\bar{h}}_{i}$
$\boldsymbol{R}_{i}$ $=\nabla_{u}\boldsymbol{\bar{h}}_{i}^{\top}\nabla_{\bar{h}\bar{h}}\mathcal{B}_{i}\nabla_{u}\boldsymbol{\bar{h}}_{i}$
As a common practice, while using the MPC in closed-loop with the robot, we perform one iteration only of the algorithm described in Sec. 1 per control loop. To increase the robustness of the method, we warm-start each call of the solver with the previous prediction just shifted by one time-step as shown by [36].
Finally, to guarantee the differentiability of the solver through JAX-based automatic differentiation, we avoided the use of $while$ loop in the code. Especially in line-search, we evaluate a fixed number of alpha values in parallel.

### v2

@@ II-D Parallel Associative Scan
In this section, we will briefly recap how parallel associative scans can be used to solve the primal problem as shown for the first time by [TemporalParallelizationLQR]. Given a sequence of elements $a_{1},\dots,a_{N}$ and an associative operator $\otimes$ defined on them, a parallel associative scan can be used to evaluate the values $s_{1},\dots,s_{N}$ as:
$(s_{1},\;s_{2},\;\dots,\;s_{N})=(a_{1},\;a_{1}\otimes a_{2},\;\dots,\;a_{1}\otimes a_{2}\otimes\dots\otimes a_{N})$ (7)
in $\mathcal{O}(\log{N})$ as shown by [scan]. This is possible since $\otimes$ is an associative operation, allowing the computation to be reorganized into smaller interdependent subproblems. To make use of this property, we first define the Conditional Value Function (CVF) $V_{i\rightarrow{}j}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$ as the minimal cost related to the optimal trajectory that goes from state $\boldsymbol{x}_{i}$ at node $i$ in the horizon to state $\boldsymbol{x}_{j}$ at node $j$, with $i<j$. Given the connection between the value function of (II-B) and its CVF, $V_{i}(\boldsymbol{x}_{i})=V_{i\rightarrow N+1}(\boldsymbol{x}_{i},\boldsymbol{x}_{N+1})$, we will use the scan to evaluate $V_{i}(\boldsymbol{x}_{i})$ for $i=0,\dots N+1$. We begin by initializing the elements $a_{i}$ with the CVF $V_{i\rightarrow{i+1}}$ defined between two adjacent states $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{i+1}$. Next, we define the operator $\otimes$ for dynamic programming as:
$V_{i\rightarrow{k}}\otimes V_{k\rightarrow{j}}=\min_{\boldsymbol{x}_{k}}\{V_{i\rightarrow{k}}+V_{k\rightarrow{j}}\}$ (8)
where $\boldsymbol{x}_{k}$ is a third intermediate state between $i$ and $j$. Note that this operator is associative since the $\min$ operator is associative. We can now perform an associative scan on the CVF to obtain $V_{i\rightarrow N+1}$ in a single pass, thus $V_{i}(\boldsymbol{x}_{i})$ for all $i$ in the horizon. To derive the combination rule for (8), we start by noting that the CVF of (II-B) is also a quadratic program with affine equality constraint (1b), which can be represented in its dual form [convex_optimization] by introducing the Lagrangian multiplier $\boldsymbol{\eta}$:
@@
where the tilde $\tilde{\cdot}$ is used to distinguish the values related to the conditional value function from the one in (4), while the subscripts are used to specify the initial and final state considered. Can be proved that given $V_{i\rightarrow{k}}$ and $V_{k\rightarrow{j}}$ that are in the same form of (9), $V_{i\rightarrow{j}}$ will still be in the form of (9) and is characterized by such combination rule:
@@
We now initialize the values $a_{i}$ of a reverse associative scan, using the conditional value function $V_{i\rightarrow{i+1}}$ characterized by:
$\boldsymbol{\tilde{A}}_{i,i+1}$ $=\boldsymbol{A_{i}}-\boldsymbol{B_{i}}\boldsymbol{R}_{i}^{-1}S_{i}^{\top},$ $\boldsymbol{\tilde{P}}_{i,i+1}$ $=\boldsymbol{Q_{i}}-\boldsymbol{S_{i}}\boldsymbol{R}_{i}^{-1}\boldsymbol{S}_{i}^{\top}$ (11)
$\boldsymbol{\tilde{C}}_{i,i+1}$ $=\boldsymbol{B_{i}}\boldsymbol{R_{i}}^{-1}\boldsymbol{B}_{i}^{\top},$ $\boldsymbol{\tilde{p}}_{i,i+1}$ $=\boldsymbol{q_{i}}-\boldsymbol{S_{i}}\boldsymbol{R}_{i}^{-1}r_{i}$
@@
$\boldsymbol{\tilde{P}}_{N,N+1}$ $=\boldsymbol{Q}_{N+1},$ $\boldsymbol{\tilde{A}}_{N,N+1}$ $=0,$ $\boldsymbol{\tilde{b}}_{N,N+1}=0$ (12)
$\boldsymbol{\tilde{p}}_{N,N+1}$ $=\boldsymbol{q}_{N+1}$ $\boldsymbol{\tilde{C}}_{N,N+1}$ $=0$
From the solution of the parallel associative scan using the combination rule in (10), we obtain $V_{i\rightarrow N+1}(\boldsymbol{x}_{i},\boldsymbol{x}_{N+1})$ for $i=0,\dots,N+1$. Finally, recalling that $V_{i\rightarrow{N+1}}(\boldsymbol{x}_{i},\boldsymbol{x}_{N+1})$ is equal to the value function $V_{i}(x_{i})$ of the problem in (II-B) and that the value function is also in the same form of (9) we get that $\boldsymbol{P}_{i}=\boldsymbol{\tilde{P}}_{i,N+1}$ and $\boldsymbol{\tilde{p}}_{i}=\boldsymbol{p}_{i,N+1}$. This means that we obtained all $\boldsymbol{P}_{i}$ and $\boldsymbol{p}_{i}$ for $i=0,\,\dots\,,N+1$ that can then be used in (II-C) to calculate $\boldsymbol{K}$ and $\boldsymbol{k}$ in on pass in parallel. Thanks to a GPU-accelerated scan implementation, we can obtain the control policy over the entire horizon with an overall computational complexity of $\mathcal{O}(\log^{2}(n)\log{N}+\log^{2}(m))$. This follows from the fact that both computing and initializing the scan involve a matrix inversion, which can be performed in parallel with complexity $\mathcal{O}(\log^{2}(n))$ [parallel_matrix_inv]. The parallel associative scan can also be used to retrieve the optimal trajectory in $\mathcal{O}(\log{n}\log{N}+\log{m})$. We start plugging into the linearized dynamics the optimal control law we evaluated, obtaining:
$\boldsymbol{\delta x}_{i+1}=\boldsymbol{\bar{A}}_{i}\boldsymbol{\delta x}_{i}+\boldsymbol{\bar{b}}_{i}$ (13)
where: $\boldsymbol{\bar{A}}_{i}=\boldsymbol{A}_{i}+\boldsymbol{B}_{i}\boldsymbol{K}_{i}$ and $\boldsymbol{\bar{b}}_{i}=\boldsymbol{B}_{i}\boldsymbol{k}_{i}+\boldsymbol{b}_{i}$ We can define the conditional optimal trajectory as follows:
$\boldsymbol{\bar{h}}_{i\rightarrow{j}}(\boldsymbol{\delta x}_{i},\boldsymbol{\delta x}_{j})=\boldsymbol{\bar{A}}_{i,j}\boldsymbol{\delta x}_{i}+\boldsymbol{\bar{b}}_{i,j}$ (14)
[TemporalParallelizationLQR] showed that from the combination of two conditional optimal trajectory $\boldsymbol{h}_{i\rightarrow{k}}$ and $\boldsymbol{h}_{k\rightarrow{j}}$ we obtain $\boldsymbol{h}_{i\rightarrow{j}}$ that can be written in the same form as (14) with:
$\boldsymbol{\bar{A}}_{i,j}=\boldsymbol{\bar{A}}_{i,k}\boldsymbol{\bar{A}}_{k,j}\quad\boldsymbol{\bar{b}}_{i,j}=\boldsymbol{\bar{A}}_{k,j}\boldsymbol{\bar{b}}_{i,k}+\boldsymbol{\bar{b}}_{k,j}$ (15)
@@
$\boldsymbol{\bar{A}}_{i,i+1}=\boldsymbol{\bar{A}}_{i}\quad\boldsymbol{\bar{b}}_{i,i+1}=\boldsymbol{\bar{b}}_{i}\\$ (16)
@@ II-E Parallel Line Search
$\boldsymbol{x}^{k+1}=\boldsymbol{x}^{k}+\alpha\boldsymbol{\delta x}$ $\boldsymbol{\lambda}^{k+1}=\boldsymbol{\lambda}^{k}+\alpha\boldsymbol{\delta\lambda}$ (18)
@@
The SQP linear rollout allows for the computation to be easily performed in parallel for the horizon. To evaluate $\alpha$, we implemented a filter line search similar to the one implemented in [ipopt]. The filter line search ensures that the step taken in the iteration reduces the constraint violation or the cost. Such a method has already been successfully deployed for legged robot control in [Perceptive_based_MPC]. We measure the satisfaction of the dynamics constraints as:
@@
As for [Perceptive_based_MPC], when the constraint violation $\theta$ is higher than a threshold, we reject the step if it further increases $\theta$. If the violation is below the threshold and the current step is a descent direction, then we apply an Armijo condition [armijo]. On the other hand, if the current iterate is not in a descent direction, we require that at least the cost or $\theta$ is decreased to accept the step. To leverage GPU parallelism, we evaluate the line-search in parallel (at a fixed grid of ten step sizes, i.e. $\alpha\in\{2^{0},...,2^{-9}\}$) and select the largest $\alpha$ that satisfies one of the line-search conditions.
@@ II-F Practical Implementation
$\boldsymbol{Q}_{N+1}$ $=\nabla_{x}\boldsymbol{\epsilon}_{N+1}^{\top}W_{T}\nabla_{x}\boldsymbol{\epsilon}_{N+1}$ $\boldsymbol{S}_{i}=\nabla_{x}\boldsymbol{\epsilon}_{i}^{\top}W_{i}\nabla_{u}\boldsymbol{\epsilon}_{i}$ (20)
$\boldsymbol{Q}_{i}$ $=\nabla_{x}\boldsymbol{\epsilon}_{i}^{\top}W_{i}\nabla_{x}\boldsymbol{\epsilon}_{i}$ $\boldsymbol{R}_{i}=\nabla_{u}\boldsymbol{\epsilon}_{i}^{\top}W_{i}\nabla_{u}\boldsymbol{\epsilon}_{i}$
@@
$\mathcal{B}({\color[rgb]{0,0,0}\xi},\mu,\delta)\begin{cases}-\mu\ln({\color[rgb]{0,0,0}\xi}),&\text{if }{\color[rgb]{0,0,0}\xi}\geq\delta,\\ \frac{\mu}{2}\left(\left(\frac{{\color[rgb]{0,0,0}\xi}-2\delta}{\delta}\right)^{2}-1\right)-\mu\ln(\delta),&\text{if }{\color[rgb]{0,0,0}\xi}<\delta.\end{cases}$ (21)
where ${\color[rgb]{0,0,0}\xi}(\boldsymbol{x},\boldsymbol{u})<0$ is the considered constraint, and $\mu$ and $\delta$ are parameters tuned for each constraint. Following the example of [Perceptive_based_MPC], to maintain the convexity of II-B the hessian approximation of the constraints is constructed as
$\boldsymbol{Q}_{N+1}$ $={\color[rgb]{0,0,0}\nabla_{x}\boldsymbol{\epsilon}_{N+1}^{\top}W_{T}\nabla_{x}\boldsymbol{\epsilon}_{N+1}}+\nabla_{x}\boldsymbol{{\color[rgb]{0,0,0}\xi}}_{N+1}^{\top}\nabla_{{\color[rgb]{0,0,0}\xi}{\color[rgb]{0,0,0}\xi}}\mathcal{B}_{T}\nabla_{x}\boldsymbol{{\color[rgb]{0,0,0}\xi}}_{N+1}$ (22)
$\boldsymbol{Q}_{i}$ $={\color[rgb]{0,0,0}\nabla_{x}\boldsymbol{\epsilon}_{i}^{\top}W_{i}\nabla_{x}\boldsymbol{\epsilon}_{i}}+\nabla_{x}\boldsymbol{{\color[rgb]{0,0,0}\xi}}_{i}^{\top}\nabla_{{\color[rgb]{0,0,0}\xi}{\color[rgb]{0,0,0}\xi}}\mathcal{B}_{i}\nabla_{x}\boldsymbol{{\color[rgb]{0,0,0}\xi}}_{i}$
$\boldsymbol{S}_{i}$ $={\color[rgb]{0,0,0}\nabla_{x}\boldsymbol{\epsilon}_{i}^{\top}W_{i}\nabla_{x}\boldsymbol{\epsilon}_{i}}+\nabla_{x}\boldsymbol{{\color[rgb]{0,0,0}\xi}}_{i}^{\top}\nabla_{{\color[rgb]{0,0,0}\xi}{\color[rgb]{0,0,0}\xi}}\mathcal{B}_{i}\nabla_{u}\boldsymbol{{\color[rgb]{0,0,0}\xi}}_{i}$
$\boldsymbol{R}_{i}$ $={\color[rgb]{0,0,0}\nabla_{u}\boldsymbol{\epsilon}_{i}^{\top}W_{i}\nabla_{u}\boldsymbol{\epsilon}_{i}}+\nabla_{u}\boldsymbol{{\color[rgb]{0,0,0}\xi}}_{i}^{\top}\nabla_{{\color[rgb]{0,0,0}\xi}{\color[rgb]{0,0,0}\xi}}\mathcal{B}_{i}\nabla_{u}\boldsymbol{{\color[rgb]{0,0,0}\xi}}_{i}$
As a common practice, while using the MPC in closed-loop with the robot, we perform one iteration only of the algorithm described in Sec. 1 per control loop. To increase the robustness of the method, we warm-start each call of the solver with the previous prediction just shifted by one time-step as shown by [rti].
Finally, to guarantee the differentiability of the solver through JAX-based automatic differentiation, we avoided the use of whilewhile loop in the code. Especially in line-search, we evaluate a fixed number of alpha values in parallel.